OpenStudy (anonymous):

Given the normal distribution with mean 100 and standard deviation 10, find the number of standard deviations the measurement is from the mean. Express the answer as a positive number in decimal form. Measurement = 139

OpenStudy (anonymous):

(measurement - mean)/SD = (139-100)/10 = 3.9 <--------